# Shuffle Products, Degenerate Affine Hecke Algebras, and Quantum Toda Lattice

### v2

@@ 5.3. Mirabolic reduction and Feigin-Odesskii shuffle algebras
$\mathfrak{m}_{N}=\mathrm{span}\left(E_{ij}\Big{|}\begin{array}[]{c}i=1\ldots N-1,\\ j=1\ldots N\end{array}\right).$
@@
(5.9) $\mathrm{Rep}(\mathrm{GL}_{N})\rightarrow\mathrm{Mod}_{\mathbf{C}[\hbar]},\qquad V\mapsto\mathrm{U}_{\hbar}(\mathfrak{m}_{N})\otimes V/\!\!/_{\psi}\,N_{+}:=(\mathrm{U}_{\hbar}(\mathfrak{m}_{N})\otimes V/\mathfrak{n}_{+}^{\psi})^{N_{+}},$
a mirabolic version of the Kostant-Whittaker reduction from ˜5.1 for free Harish-Chandra bimodules. It follows from [Kal25, Theorem 5.3] that the action of the Kirillov projector of Theorem˜5.2 is well-defined on $\mathrm{U}_{\hbar}(\mathfrak{m}_{N})\otimes V/\mathfrak{n}_{+}^{\psi}$ and gives an isomorphism
(5.10) $V[\hbar]\xrightarrow{\sim}\mathrm{U}_{\hbar}(\mathfrak{m}_{N})\otimes V/\!\!/_{\psi}\,N_{+},\qquad v\mapsto P_{\psi}vP_{\psi}.$
The following is an immediate corollary of ˜5.4 and Proposition˜5.10.
@@
    $\mathbf{C}[\hbar][\Omega_{1},\ldots,\Omega_{k}]_{<N}\xrightarrow{\sim}\mathrm{U}_{\hbar}(\mathfrak{m}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}/\!\!/_{\psi}\,N_{+},\qquad f(\vec{\Omega})\mapsto f(\vec{\Omega})(v_{1}^{\otimes k}).$
  2. (2)
@@
    $\varphi\colon(\mathbf{C}^{N})^{\otimes k}[\hbar]\xrightarrow{\sim}\mathbf{C}[\hbar][\Omega_{1},\ldots,\Omega_{k}]_{<N}$
@@
    $\varphi(v_{i_{1}}\otimes\ldots\otimes v_{i_{k}})$ $=\psi^{k-i_{1}-\ldots-i_{k}}\Omega_{1}^{i_{1}-1}\ldots\Omega_{k}^{i_{k}-1},$
    $\varphi(v_{i_{\sigma(1)}}\otimes\ldots\otimes v_{i_{\sigma(k)}})$ $=\sigma^{\hbar}\cdot\varphi(v_{i_{1}}\otimes\ldots\otimes v_{i_{k}}),$
@@
$v_{i_{1}}\cdot\ldots\cdot v_{i_{k}}:=\frac{1}{k!}\sum_{\sigma\in S_{k}}v_{i_{\sigma(1)}}\otimes\ldots\otimes v_{i_{\sigma(k)}},\qquad N\geq i_{1}\geq\ldots\geq i_{k}\geq 1.$
Therefore, by Proposition˜5.11 and Proposition˜2.7, we have
$P_{\psi}v_{i_{1}}\cdot\ldots\cdot v_{i_{k}}P_{\psi}\mapsto\frac{\psi^{k-i_{1}-\ldots-i_{k}}}{k!}\mathrm{Sym}_{k}\left(\prod_{i<j}\frac{\Omega_{i}-\Omega_{j}+\hbar}{\Omega_{i}-\Omega_{j}}\cdot\Omega_{1}^{i_{1}-1}\ldots\Omega_{k}^{i_{k}-1}\right),$
which is a rational analog of the Hall-Littlewood polynomial associated to the partition $(i_{1}-1,\ldots,i_{k}-1)$.
@@
Consider $\Lambda^{k}\mathbf{C}^{N}\subset(\mathbf{C}^{N})^{\otimes k}$. It has a basis
$v_{i_{1}}\wedge\ldots\wedge v_{i_{k}}:=\frac{1}{k!}\sum_{\sigma\in S_{k}}(-1)^{\sigma}v_{i_{\sigma(1)}}\otimes\ldots\otimes v_{i_{\sigma(k)}},\qquad N\geq i_{1}>\ldots>i_{k}\geq 1.$
Therefore, by Proposition˜5.11 and Proposition˜2.7, we have
$P_{\psi}v_{i_{1}}\wedge\ldots\wedge v_{i_{k}}P_{\psi}\mapsto\frac{\psi^{k-i_{1}-\ldots-i_{k}}}{k!}\prod_{i<j}(\Omega_{i}-\Omega_{j}-\hbar)\cdot s_{\lambda}(\vec{\Omega}),\qquad\lambda=(i_{1}-k,\ldots,i_{k}-1),$
where $s_{\lambda}(\vec{\Omega})$ is the Schur polynomial associated to partition $\lambda$, see [Mac95, Section I.3], and the factor can be viewed as a deformation of the Vandermonde determinant.
By [Kal25, Theorem 6.3], the functor (5.9) is monoidal, in particular, if $A$ is an algebra in $\mathrm{Rep}(\mathrm{GL}_{N})$, one can define a mirabolic analog $\mathrm{U}_{\hbar}(\mathfrak{m}_{N})\ltimes A$ of ˜3.4, and its reduction $\mathrm{U}_{\hbar}(\mathfrak{m}_{N})\ltimes A/\!\!/_{\psi}\,N_{+}$ is an algebra as well, similarly to ˜5.6. Recall the rational Feigin-Odesskii shuffle algebra $^{N}\mathcal{S}^{\hbar}$ of ˜2.9. The following is a corollary of Theorem˜2.11, Theorem˜5.9, and Example˜5.12.
@@
As we mentioned in the introduction, the shuffle algebra $^{N}\mathcal{S}^{\hbar}$ is a rational degeneration of its elliptic version, in view of Section˜2 and the results of [EH04]. The original construction of [OF89] of the latter used Belavin’s $R$-matrix that can be associated to a stable bundle of rank $N$ and degree 1 (for instance, see “Geometric interpretation” after [LOZ15, (1.56)] in loc. cit. for precise statement). Unfortunately, we do not know an interpretation of Theorem˜5.14 in these terms. However, it seems to be a step in the right direction based on the following observation.
Let $E$ be an elliptic curve with the origin $0\in E$. Denote by $\xi_{n,r}$ a unique stable bundle $E$ of rank $r$ whose determinant is $\mathcal{O}_{E}(r\cdot 0)$. As it is argued in [FO98], the elliptic shuffle algebra quantizes a certain Poisson bracket on the space of extensions $\mathrm{Ext}^{1}(\xi_{N,1},\xi_{0,1})$. Its ring of functions is the symmetric algebra on
@@
Following loc. cit., we can apply the Fourier-Mukai transform
$\Phi(\xi_{-N,1})=\xi_{1,N}[-1],\qquad\Phi(\xi_{0,1})[-1]=\mathcal{O}_{0}[-1]$
where $\mathcal{O}_{0}$ is the skyscraper sheaf at the origin. Therefore, it transforms $\mathrm{Hom}^{0}(\xi_{-N,1},\xi_{0,1})$ to the space of morphisms $\xi_{1,N}\rightarrow\mathcal{O}_{0}$ which can be interpreted as a mirabolic reduction of the fiber of $\xi_{1,N}$ at the origin. To relate to the Kostant-Whittaker reduction, recall that it quantizes the Kostant slice [Kos78]. The latter can be identified with $\mathfrak{h}_{N}/W$. Its elliptic version is the coarse moduli space of degree 0 semistable bundles of rank $N$. At the same time, we can complete a non-zero $\varphi\in\mathrm{Hom}^{0}(\xi_{-N,1},\xi_{0,1})$ to an exact triangle
@@
6. Finite Miura transform
In this section, we relate the Kostant-Whittaker reduction of Section˜5 to the parabolic restriction of Section˜4 following the construction of [GK22, Section 6].
@@ 6. Finite Miura transform
Recall the Harish-Chandra map of Example˜4.4:
(6.1) $\mathbf{Z}_{\mathfrak{gl}_{N}}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}/\!\!/N_{-}\cong\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}},\qquad A(u)\mapsto PA(u)=\prod_{i=1}^{N}(u-w_{i}).$
Following [GK22, Section 6], consider the map
@@
$p_{X}\colon X/\!\!/_{\psi}\,N_{+}=(X/\mathfrak{n}_{+}^{\psi})^{N_{+}}\rightarrow\mathfrak{n}_{-}\backslash X/\mathfrak{n}_{+}^{\psi}$
@@
(6.3) $X^{\mathrm{gen}}/\!\!/N_{-}=(\mathfrak{n}_{-}\backslash X^{\mathrm{gen}})^{N_{-}}\rightarrow\mathfrak{n}_{-}\backslash X^{\mathrm{gen}}/\mathfrak{n}^{\psi}_{+}.$
@@
  1. (1)
    [GK22, Lemma 6.2.1] The map (6.2) is an isomorphism.
  2. (2)
    [Kal25, Proposition 7.6] The map (6.3) is an isomorphism.
@@
For free Harish-Chandra bimodules, the isomorphism of the second part holds without genericity assumption by [GK22, Theorem 7.1.8].
@@
In view of the map (6.3), the Miura transform induces a $W$-action on $X^{\mathrm{gen}}/\!\!/N_{-}$ extending the one of Example˜4.4 such that $X/\!\!/_{\psi}\,N_{+}\subset(X^{\mathrm{gen}}/\!\!/N_{-})^{W}$.
@@
Consider $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}$. Combining Theorem˜5.9, Theorem˜4.11, and ˜6.6, we obtain an algebra homomorphism
(6.4) $\mathrm{Y}_{-N,N}(\mathfrak{gl}_{2})^{\geq 0}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/N_{-}\cong{}^{N}\tilde{\mathcal{A}}_{\hbar}^{-}$
@@
The map (6.4) coincides with the GKLO map of Theorem˜2.16.
@@
It follows from Example˜4.4 that
@@
By arguments of [Kal25, Corollary 7.12] adapted to the positive case, we have
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}\ni\frac{1}{u-\Omega_{1}}(v_{1})\mapsto\sum_{i=1}^{N}\frac{1}{u-w_{i}}\prod_{j<i}(w_{i}-w_{j})^{-1}\cdot Pv_{i}P\in\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\otimes\mathbf{C}^{N}/\!\!/N_{-}.$
@@
7. Toda lattice
In this section, we show an isomorphism between the quantum Toda lattice for $\mathrm{GL}_{N}$ and the truncated shifted Yangian from Section˜2. The main object of this section is the algebra $\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$ of differential operators from Section˜1.3. Recall that it has two $\mathrm{GL}_{N}$ actions.
Notation 7.1.
In what follows, we use the superscript $R$ (resp. $L$) to denote everything related to the right action (resp. left action) of $\mathrm{GL}_{N}$ on itself. For instance, let $\mathrm{HC}_{\hbar}(\mathrm{GL}_{N}^{R})$ be the Harish-Chandra category $\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})$ such that $\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$ is an object $\mathrm{HC}_{\hbar}(\mathrm{GL}_{N}^{R})$ via:

### v4

@@ 6.3 Mirabolic reduction and Feigin–Odesskii shuffle algebras
$\mathfrak{m}_{N}=\mathrm{span}(E_{ij}\mid i=1,\dots,N-1,\,j=1,\dots,N).$
@@
$\mathrm{Rep}(\mathrm{GL}_{N})\rightarrow\mathrm{Mod}_{\mathbf{C}[\hbar]},\qquad V\mapsto\mathrm{U}_{\hbar}(\mathfrak{m}_{N})\otimes V/\!\!/_{\psi}\,N_{+}:=\bigl(\mathrm{U}_{\hbar}(\mathfrak{m}_{N})\otimes V/\mathfrak{n}_{+}^{\psi}\bigr)^{N_{+}},$ (6.9)
a mirabolic version of the Kostant–Whittaker reduction from Definition 6.1 for free Harish-Chandra bimodules. It follows from [27, Theorem 5.3] that the action of the Kirillov projector of (6.2) is well-defined on $\mathrm{U}_{\hbar}(\mathfrak{m}_{N})\otimes V/\mathfrak{n}_{+}^{\psi}$ and gives an isomorphism
$V[\hbar]\xrightarrow{\sim}\mathrm{U}_{\hbar}(\mathfrak{m}_{N})\otimes V/\!\!/_{\psi}\,N_{+},\qquad v\mapsto P_{\psi}vP_{\psi}.$ (6.10)
The following is an immediate corollary of Lemma 6.4 and Proposition 6.10.
@@
    $\mathbf{C}[\hbar][\Omega_{1},\dots,\Omega_{k}]_{<N}\xrightarrow{\sim}\mathrm{U}_{\hbar}(\mathfrak{m}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}/\!\!/_{\psi}\,N_{+},\qquad f\bigl(\vec{\Omega}\bigr)\mapsto f\bigl(\vec{\Omega}\bigr)\bigl(v_{1}^{\otimes k}\bigr).$
  2. $2.$
@@
    $\varphi\colon\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}[\hbar]\xrightarrow{\sim}\mathbf{C}[\hbar][\Omega_{1},\dots,\Omega_{k}]_{<N}$
@@
    $\varphi(v_{i_{1}}\otimes\dots\otimes v_{i_{k}})=\psi^{k-i_{1}-\dots-i_{k}}\Omega_{1}^{i_{1}-1}\cdots\Omega_{k}^{i_{k}-1},$
    $\varphi\bigl(v_{i_{\sigma(1)}}\otimes\dots\otimes v_{i_{\sigma(k)}}\bigr)=\sigma^{\hbar}\cdot\varphi(v_{i_{1}}\otimes\dots\otimes v_{i_{k}}),$
@@
$v_{i_{1}}\cdots v_{i_{k}}:=\frac{1}{k!}\sum_{\sigma\in S_{k}}v_{i_{\sigma(1)}}\otimes\dots\otimes v_{i_{\sigma(k)}},\qquad N\geq i_{1}\geq\dots\geq i_{k}\geq 1.$
Therefore, by Propositions 6.11 and 3.7, we have
$P_{\psi}v_{i_{1}}\cdots v_{i_{k}}P_{\psi}\mapsto\frac{\psi^{k-i_{1}-\dots-i_{k}}}{k!}\mathrm{Sym}_{k}\biggl(\prod_{i<j}\frac{\Omega_{i}-\Omega_{j}+\hbar}{\Omega_{i}-\Omega_{j}}\cdot\Omega_{1}^{i_{1}-1}\cdots\Omega_{k}^{i_{k}-1}\biggr),$
which is a rational analog of the Hall–Littlewood polynomial associated to the partition $({i_{1}-1},\allowbreak\dots,{i_{k}-1})$.
@@
Consider $\Lambda^{k}\mathbf{C}^{N}\subset\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}$. It has a basis
$v_{i_{1}}\wedge\dots\wedge v_{i_{k}}:=\frac{1}{k!}\sum_{\sigma\in S_{k}}(-1)^{\sigma}v_{i_{\sigma(1)}}\otimes\dots\otimes v_{i_{\sigma(k)}},\qquad N\geq i_{1}>\dots>i_{k}\geq 1.$
Therefore, by Proposition 6.11 and Proposition 3.7, we have
$P_{\psi}v_{i_{1}}\wedge\dots\wedge v_{i_{k}}P_{\psi}\mapsto\frac{\psi^{k-i_{1}-\dots-i_{k}}}{k!}\prod_{i<j}(\Omega_{i}-\Omega_{j}-\hbar)\cdot s_{\lambda}\bigl(\vec{\Omega}\bigr),\qquad\lambda=(i_{1}-k,\dots,i_{k}-1),$
where $s_{\lambda}\bigl(\vec{\Omega}\bigr)$ is the Schur polynomial associated to partition $\lambda$, see [36, Section I.3], and the factor can be viewed as a deformation of the Vandermonde determinant.
By [27, Theorem 6.3], the functor (6.9) is monoidal, in particular, if $A$ is an algebra in $\mathrm{Rep}(\mathrm{GL}_{N})$, one can define a mirabolic analog $\mathrm{U}_{\hbar}(\mathfrak{m}_{N})\ltimes A$ of Definition 4.4, and its reduction $\mathrm{U}_{\hbar}(\mathfrak{m}_{N})\ltimes A/\!\!/_{\psi}\,N_{+}$ is an algebra as well, similarly to Definition 6.6. Recall the rational Feigin–Odesskii shuffle algebra $^{N}\!\mathcal{S}^{\hbar}$ of Definition 3.9. The following is a corollary of Theorem 3.11, Theorem 6.9, and Example 6.12.
@@
As we mentioned in the introduction, the shuffle algebra $^{N}\!\mathcal{S}^{\hbar}$ is a rational degeneration of its elliptic version, in view of Section 3 and the results of [10]. The original construction of [40] of the latter used Belavin’s $R$-matrix that can be associated to a stable bundle of rank $N$ and degree 1 (for instance, see “Geometric interpretation” after [34, formula (1.56)] in loc. cit. for precise statement). Unfortunately, we do not know an interpretation of Theorem 6.14 in these terms. However, it seems to be a step in the right direction based on the following observation.
Let $E$ be an elliptic curve with the origin $0\in E$. Denote by $\xi_{n,r}$ a unique stable bundle $E$ of rank $r$ whose determinant is $\mathcal{O}_{E}(n\cdot 0)$. As it is argued in [12], the elliptic shuffle algebra quantizes a certain Poisson bracket on the space of extensions $\mathrm{Ext}^{1}(\xi_{N,1},\xi_{0,1})$. Its ring of functions is the symmetric algebra on
@@
Following loc. cit., we can apply the Fourier–Mukai transform
$\Phi(\xi_{-N,1})=\xi_{1,N}[-1],\qquad\Phi(\xi_{0,1})[-1]=\mathcal{O}_{0}[-1],$
where $\mathcal{O}_{0}$ is the skyscraper sheaf at the origin. Therefore, it transforms $\mathrm{Hom}^{0}(\xi_{-N,1},\xi_{0,1})$ to the space of morphisms $\xi_{1,N}\rightarrow\mathcal{O}_{0}$ which can be interpreted as a mirabolic reduction of the fiber of $\xi_{1,N}$ at the origin. To relate to the Kostant–Whittaker reduction, recall that it quantizes the Kostant slice [33]. The latter can be identified with $\mathfrak{h}_{N}/W$. Its elliptic version is the coarse moduli space of degree 0 semistable bundles of rank $N$. At the same time, we can complete a non-zero $\varphi\in\mathrm{Hom}^{0}(\xi_{-N,1},\xi_{0,1})$ to an exact triangle
@@
7 Finite Miura transform
In this section, we relate the Kostant–Whittaker reduction of Section 6 to the parabolic restriction of Section 5 following the construction of [22, Section 6], a finite analog of the Miura transform for W-algebras. We compare it with the GKLO homomorphism for the positive part $\mathrm{Y}_{-N,N}(\mathfrak{sl}_{2})^{\geq 0}$ from Section 3 in Theorem 7.7. It will be extended to the whole Yangian in Section 8 and used to prove the main result Theorem 8.10.
In this section, we relate the Kostant–Whittaker reduction of Section 6 to the parabolic restriction of Section 5 following the construction of [22, Section 6].
@@ 7 Finite Miura transform
Recall the Harish-Chandra map of Example 5.4:
$\mathbf{Z}_{\mathfrak{gl}_{N}}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}/\!\!/N_{-}\cong\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}},\qquad A(u)\mapsto PA(u)=\prod_{i=1}^{N}(u-w_{i}).$ (7.1)
Following [22, Section 6], consider the map
@@
$p_{X}\colon\ X/\!\!/_{\psi}\,N_{+}=\bigl(X/\mathfrak{n}_{+}^{\psi}\bigr)^{N_{+}}\rightarrow\mathfrak{n}_{-}\backslash X/\mathfrak{n}_{+}^{\psi}$
@@
$X^{\mathrm{gen}}/\!\!/N_{-}=(\mathfrak{n}_{-}\backslash X^{\mathrm{gen}})^{N_{-}}\rightarrow\mathfrak{n}_{-}\backslash X^{\mathrm{gen}}/\mathfrak{n}^{\psi}_{+}.$ (7.3)
@@
  1. $1.$
    The map (7.2) is an isomorphism [22, Lemma 6.2.1].
  2. $2.$
    The map (7.3) is an isomorphism [27, Proposition 7.6].
@@
For free Harish-Chandra bimodules, the isomorphism of the second part holds without genericity assumption by [22, Theorem 7.1.8].
@@
In view of the map (7.3), the Miura transform induces a $W$-action on $X^{\mathrm{gen}}/\!\!/N_{-}$ extending the one of Example 5.4 such that $X/\!\!/_{\psi}\,N_{+}\subset(X^{\mathrm{gen}}/\!\!/N_{-})^{W}$.
@@
Consider $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}$. Combining Theorems 6.9 and 5.11, and Lemma 7.6, we obtain an algebra homomorphism
$\mathrm{Y}_{-N,N}(\mathfrak{gl}_{2})^{\geq 0}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/N_{-}\cong{}^{N}\!\tilde{\mathcal{A}}_{\hbar}^{-}.$ (7.4)
@@
The map (7.4) coincides with the GKLO map of Theorem 3.16.
@@
It follows from Example 5.4 that
@@
By arguments of [27, Corollary 7.12] adapted to the positive case, we have
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}\ni\frac{1}{u-\Omega_{1}}(v_{1})$
$\qquad{}\mapsto\sum_{i=1}^{N}\frac{1}{u-w_{i}}\prod_{j<i}(w_{i}-w_{j})^{-1}\cdot Pv_{i}P\in\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\otimes\mathbf{C}^{N}/\!\!/N_{-}.$
@@
8 Toda lattice
In this section, we show an isomorphism between the quantum Toda lattice for $\mathrm{GL}_{N}$ and the truncated shifted Yangian from Section 3 in Theorem 8.10. The result is based on identification of the Miura transform of the quantum Toda lattice with the algebra of difference operators, see Theorem 8.8, and comparing their images inside the latter. In Section 8.1, we compare our result with its geometric counterpart, in terms of equivariant homology of the affine Grassmannian. Finally, in Section 8.2, we show that the images of matrix coefficient functions associated to fundamental weights are given by degeneration of (rational) Macdonald operators.
The main object of this section is the algebra $\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$ of differential operators from Section 2.3. Recall that it has two $\mathrm{GL}_{N}$ actions.
Notation 8.1.
In what follows, we use the superscript $R$ (resp. $L$) to denote everything related to the right action (resp. left action) of $\mathrm{GL}_{N}$ on itself. For instance, let $\mathrm{HC}_{\hbar}\bigl(\mathrm{GL}_{N}^{R}\bigr)$ be the Harish-Chandra category $\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})$ such that $\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$ is an object $\mathrm{HC}_{\hbar}\bigl(\mathrm{GL}_{N}^{R}\bigr)$ via: